OpenStudy (anonymous):

what is the slope lines through points (6,3) and (14,19)

OpenStudy (anonymous):

HINT: y - y1 = (y2-y1)/(x2-x1) (x-x1)

OpenStudy (anonymous):

\[y2-y1/ x2-x1\] = slope

OpenStudy (anonymous):

oooo ok thanks

OpenStudy (anonymous):

2

OpenStudy (anonymous):

what

OpenStudy (anonymous):

\[\frac{ 19-3 }{ 14-6 }=\frac{ 16 }{ 8 }=2\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(: